# 3Doodle: Compact Abstraction of Objects with 3D Strokes

### v1

@@ Abstract
While free-hand sketching has long served as an efficient representation to convey characteristics of an object, they are often subjective, deviating significantly from realistic representations. Moreover, sketches are not consistent for arbitrary viewpoints, making it hard to catch 3D shapes. We propose 3Dooole, generating descriptive and view-consistent sketch images given multi-view images of the target object. Our method is based on the idea that a set of 3D strokes can efficiently represent 3D structural information and render view-consistent 2D sketches. We express 2D sketches as a union of view-independent and view-dependent components. 3D cubic Bézier curves indicate view-independent 3D feature lines, while contours of superquadrics express a smooth outline of the volume of varying viewpoints. Our pipeline directly optimizes the parameters of 3D stroke primitives to minimize perceptual losses in a fully differentiable manner. The resulting sparse set of 3D strokes can be rendered as abstract sketches containing essential 3D characteristic shapes of various objects. We demonstrate that 3Doodle can faithfully express concepts of the original images compared with recent sketch generation approaches.
@@
Figure 1: We propose 3Doodle, a method to draw 3D strokes from multi-view images. Our approach can express the target object with a highly compact set of geometric primitives, and the recovered 3D strokes are rendered to generate view-coherent sketches. We show the results obtained from 3Doodle.
1 Introduction
@@
Free-hand sketching is an effective tool for visual communication, as only a few sketch lines can summarize perceptually distinctive components. Sketch strokes are also intuitive tools to interact with users in editing an image in many graphics applications [44, 46]. Previous works further explore the connection between 3D structure and conventional sketch lines and show promising results in converting sketches into 3D objects [21, 22, 13, 49] or extracting sketch lines from detailed 3D geometry [25, 5, 24]. Like image editing applications, user sketches provide an interactive tool to edit the 3D models [28, 2].
@@
The aforementioned works assume full 3D models are available for generating sketches or editing 3D objects with line strokes. However, it is not trivial to formulate the exact mapping from artistic sketch curves on-screen space to the underlying 3D structure of the object. While sketches may contain important spatial arrangements of the object, hand-drawn sketches do not strictly abide by the 3D geometry, either due to a lack of delicate skills or the artist’s creative expression. Therefore, we cannot extract consistent 3D lines from 2D sketch lines extracted individually from different viewpoints. Previous works also propose sophisticated techniques for converting 3D models, such as high-resolution triangular mesh, into sketch images, which result in delicate renderings rather than sparse and abstract strokes [9, 17]. The task does not rely on a data-driven method, as it is hard to obtain a dataset of 3D object-sketch pairs in sufficient quantity and quality.
@@
We directly optimize for proposed lightweight 3D primitives given only multi-view images of an object as input. We do not need to individually stylize the input images into sketches, and we also do not require any explicit 3D mesh or NeRF volume [29]. Most importantly, our rendering pipeline is fully differentiable for a small set of parameters representing the Bézier curves and superquadrics. We can then directly optimize the parameters to minimize the popular perceptual loss functions, such as CLIP [36] and LPIPS [48], so that the rendered images are perceived as hand-drawn 2D sketches from various viewpoints. Further, the coherent 3D structure summarizes important geometric features that satisfy multi-view consistency, which can potentially be extended to various 3D manipulation tasks.
@@
We demonstrate that the proposed compact 3D representation can easily mimic an expressive hand drawing of diverse pictured objects. Our 3Doodle faithfully expresses concepts of the target shape and provides view-consistent sketch images.
2 Related Works
@@
Generating sketches from photos of objects suffers from a large domain gap between pictures and sketches. Images contain dense, detailed evidence of real objects. In contrast, sketches are sparse and abstract, as they are traditionally a quick tool for an artist to summarize the essential components of the given object on an image plane. While extracting edges from image pixels [8, 39] can also create a line drawing of the object, they do not enjoy the compactness and implicit semantic mapping of conventional sketches. Numerous works overcome the domain gap by exploiting datasets composed of pairs of images and sketches [38, 18]. However, obtaining high-quality sketches of given images on a large scale is difficult. Most sketch generation methods only cover specific categories in datasets, such as human portrait [6, 43], or a limited number of object categories  [45, 11, 14]. A few works alleviate the necessity of paired datasets and generate sketches of arbitrary objects with style transfer  [15, 26] or image translation [52, 16]. Although they can mimic the style of the free-hand sketches to some extent, the generated pixels are still composed of dense information from original images. The images in the converted style do not necessarily maintain compact and sparse representation, which are among the most powerful features of sketches.
3Doodle proposes maintaining the sparse characteristics by directly optimizing a set of strokes, similar to recent approaches [23, 40]. The direct optimization does not require any dataset to train and, therefore, can generate sketches of arbitrary class objects. However, previous works optimize 2D strokes in the input image view without explicit 3D structure. Instead, we optimize 3D strokes with a novel differentiable rendering pipeline. 3Doodle, therefore, generates a compact 3D representation, which can be rendered in various viewpoints to deliver abstract semantic structure.
@@
Non-photorealistic rendering (NPR) renders 3D representations that exhibit various artistic styles. Among possible variations, our sketch generation task particularly relates to line drawing from 3D models. Line drawings reflect how humans understand the world [5]. Previous works on line-drawing extraction trace feature lines on surface extrema [30] in addition to view-dependent geometric boundaries such as occluding contours [5] or suggestive contours [9]. A more recent method [25] combines an entirely geometric approach with image translation and trains a neural network that generates sketches from 3D models. However, all the works rely on having an accurate mesh model, which may not be accessible in real-world scenarios. Further, geometric processing results deviate from typical free-hand drawings because they ignore the effects of texture variations, and lines are densely populated compared to rough sketches.
Recent NPR approaches propose relaxing the mesh requirement and using the 3D volume of neural radiance fields (NeRF) trained from multi-view images instead. They transform pre-trained NeRF into the style of an exemplar image to generate multi-view consistent sketches [47, 50]. While the pipelines use more casual input than a complete mesh, they need heavy computation first to train a neural volume. Neural Edge Fields (NEF) [42] also extracts a compact set of parametric curves as 3Doodle, but it is converted with additional 2D edge input from pre-trained neural fields. Furthermore, the 3D lines from NEF cannot represent view-dependent contours, which are crucial in expressing the overall shape of the object.
3 3Doodle
Figure 2: Overview of 3Doodle. We generate a compact 3D geometric representation from multi-view images of objects. We separately define view-independent stroke ($\mathcal{S}^{\text{3D}}_{\text{ind}}$) and view-dependent stroke ($\mathcal{S}^{\text{3D}}_{\text{dep}}$). We represent a view-independent stroke as a set of 3D Bézier curves and a view-dependent stroke as a contour of superquadrics. (Sec. 3.1) We also propose a fully differentiable rendering method to render sketches from the 3D strokes. (Sec. 3.2) Finally, our 3D stroke parameters are directly optimized with perceptual losses. (Sec. 3.3)
We first introduce our 3D geometric primitives that generate sketch images in Section 3.1. Then our differentiable rendering pipeline in  Section 3.2 converts the proposed primitives into an abstract sketch image. The compact set of parameters is effectively optimized to minimize the perceptual loss function as described in Section 3.3.
3.1 Sketch Representation
@@
$\mathcal{S}^{\text{3D}}=\mathcal{S}_{\text{ind}}^{\text{3D}}\cup\mathcal{S}_{\text{dep}}^{\text{3D}}.$ (1)
@@
$\mathcal{S}_{\text{ind}}^{3\text{D}}=\left\{B^{\text{3D}}(p_{i})\right\}_{i=1}^{N_{\text{ind}}},\quad p_{i}=(p_{i}^{0},p_{i}^{1},p_{i}^{2},p_{i}^{3})$ (2)
@@
$B^{\text{3D}}(t;p_{i})=\sum_{j=0}^{3}b_{j}(t)p_{i}^{j},\quad b_{j}(t)={\binom{3}{j}}t^{j}(1-t)^{3-j},$ (3)
@@
Additionally, view-dependent components $\mathcal{S}_{\text{dep}}^{3\text{D}}$ encapsulate the 3D volume of the given object. Inspired by the previous works [35, 1, 32, 33], we use the composition of superquadradics [3] as a compact parametric representation that can express various 3D shapes. More concretely,
$\mathcal{S}_{\text{dep}}^{3\text{D}}(\mathbf{d})=\sigma_{\text{contour}}\left(\bigcup_{i=1}^{N_{\text{dep}}}S(\theta_{i}),\mathbf{d}\right),$ (4)
where $N_{\text{dep}}$ is the number of superquadrics, and $\sigma_{\text{contour}}(G,\mathbf{d})$ is the view-dependent volume density function that generates the contour of any geometric volume $G$ at the given viewing direction $\mathbf{d}$. We will discuss contour extraction in Section 3.2.
@@
$f(\mathbf{x};\alpha_{i},\epsilon_{i})\,{=}\,\left(\left(\frac{x}{\alpha_{i,1}}\right)^{\frac{2}{\epsilon_{i,2}}}\!{+}\,\left(\frac{y}{\alpha_{i,2}}\right)^{\frac{2}{\epsilon_{i,2}}}\right)^{\frac{\epsilon_{i,2}}{\epsilon_{i,1}}}\!{+}\,\left(\frac{z}{\alpha_{i,3}}\right)^{\frac{2}{\epsilon_{i,1}}},$ (5)
@@
$S(\mathbf{x};\theta_{i})=f\left(\mathbf{R}_{i}^{-1}(\mathbf{x}-\mathbf{t}_{i});\alpha_{i},\epsilon_{i}\right).$ (6)
@@
$S(\mathbf{x};\theta)=\bigcup_{i=1}^{N_{\text{dep}}}S(\mathbf{x};\theta_{i})=\min\limits_{i}S(\mathbf{x};\theta_{i}).$ (7)
3.2 Differentiable Rendering of Sketch Lines
Given the 3D geometric primitives, we can draw black sketch lines on a white canvas $\mathcal{R}(\mathcal{S}^{\text{3D}})$. A complete sketch is the union of view-independent and view-dependent components. We propose a fully differentiable rendering pipeline to directly optimize the 3D parameters to produce abstract sketch images of the target objects.
@@ 3.2 Differentiable Rendering of Sketch Lines
We provide the proof of Theorem 1 in Appendix A. Following Theorem 1, we can render the 3D Bézier curve $B^{\text{3D}}$ at a given viewpoint by 1) perspective-projecting four control points and then 2) rendering a 2D Bézier curve with existing differentiable rasterizer [23]. The rendering pipeline is fully differentiable since the projection operation and rasterizer are both differentiable.
Figure 3: (a) We visualize the volume density $\sigma_{\text{vol}}$ and smooth surface volume density $\sigma_{\text{surf}}$. (b) We display the volume-rendered results of each volume density component. One can obtain the contour of geometric volume by volume rendering our proposed view-dependent contour volume density $\sigma_{\text{contour}}$.
@@
We provide the formulation to generate contour sketches that are differentiable to the superquadric parameters. Volume rendering exploits smooth transition in volume density and provides a powerful tool to propagate image-space gradients to 3D geometry, as demonstrated by neural radiance fields [29]. Similarly, we propose a differentiable pipeline by defining smooth density fields representing sketch lines. To allow continuous variation, ISCO [1] proposes volume density of superquadric
$\sigma_{\text{vol}}(\mathbf{x})=\text{sigmoid}(\gamma(1-S(\mathbf{x})),$ (8)
where $S$ is an implicit surface function defined in Equations 5, 6 and 7. $\gamma$ controls the slope of the shape boundary. For sketch generation, we revise the volume density function to be concentrated near the surface $S=1$:
$\sigma_{\text{surf}}(\mathbf{x})\,{=}\,a\cdot\text{sigmoid}\left(\frac{1}{\gamma(1-S(\mathbf{x}))^{2}\,{+}\,\epsilon}-\gamma(1-S(\mathbf{x}))^{2}\,{+}\,b\right),$ (9)
where $a,b\in\mathbb{R}^{+}$ are hyperparameters determining the intensity and thickness of the surface volume. We illustrate the $\sigma_{\text{vol}}$ and $\sigma_{\text{surf}}$ in one-dimensional example in Figure 3 (a).
@@
$\sigma_{\text{contour}}(\mathbf{x},\mathbf{d})=(1-(\mathbf{n}(\mathbf{x})\cdot\mathbf{d})^{\beta})\sigma_{\text{surf}},$ (10)
where $\mathbf{d}$ is a viewing direction and $\mathbf{n}$ is a normal direction at the querying point $\mathbf{x}$. Note that the normal direction $\mathbf{n}(\mathbf{x})=\left[\frac{\partial S}{\partial x},\frac{\partial S}{\partial y},\frac{\partial S}{\partial z}\right]$ can be calculated in a closed form with analytic derivative. Then we render the view-dependent sketch $\mathcal{R}(\mathcal{S}_{\text{dep}}^{\text{{3D}}})$ using volume rendering technique and numerical quadrature approximation [27] with our contour volume density $\sigma_{\text{contour}}$.
3.3 Optimization
Given the differentiable pipeline, we optimize the parameters of our sparse 3D primitives to generate sketch images. While there exists a significant domain gap between photo-realistic input images and the abstract sketch [40], we design a loss function to encourage a balance between structural components and semantic perception compared to input multi-view images. Specifically, we use LPIPS loss [48] to capture rough geometric layout and CLIP score [36] to maintain the high-level semantic meaning:
$\begin{split}\mathcal{L}=\sum_{I\in\mathcal{I}}\lambda~{}\rho\left(\text{LPIPS}(I,\mathcal{R}(\mathcal{S}^{\text{3D}})),\alpha,c\right)\\ +\,\text{dist}(\text{CLIP}(I),\text{CLIP}(\mathcal{R}(\mathcal{S}^{\text{3D}}))),\end{split}$ (11)
where $I$ is the image sampled from the input multi-view images $\mathcal{I}$, and $\text{dist}(a,b)=1-\frac{a\cdot b}{\|a\|\cdot\|b\|}$ is a cosine distance. $\rho(x,\alpha,c)$ is a robust loss function [4] and we use $\alpha=1,c=0.1$ for all experiments. More studies regarding the effects of loss terms can be found in the Section 4. Our sketch parameters are directly optimized with Adam optimizer [19] by minimizing the loss term in Equation 11.
Due to the highly non-convex nature of our minimizing objective, appropriate initialization leads to fast and robust optimization. We leverage the SfM (structure-from-motion) [37] point clouds obtained while estimating the camera poses from input multi-view images. We sample SfM points using farthest point sampling and initialize them to be locations of 3D Bézier curve control points and the center of superquadrics. We further discuss the effects of the initialization method in Section 4.
4 Experiments

### v2

@@ Abstract.
While free-hand sketching has long served as an efficient representation to convey characteristics of an object, they are often subjective, deviating significantly from realistic representations. Moreover, sketches are not consistent for arbitrary viewpoints, making it hard to catch 3D shapes. We propose 3Dooole, generating descriptive and view-consistent sketch images given multi-view images of the target object. Our method is based on the idea that a set of 3D strokes can efficiently represent 3D structural information and render view-consistent 2D sketches. We express 2D sketches as a union of view-independent and view-dependent components. 3D cubic Bézier curves indicate view-independent 3D feature lines, while contours of superquadrics express a smooth outline of the volume of varying viewpoints. Our pipeline directly optimizes the parameters of 3D stroke primitives to minimize perceptual losses in a fully differentiable manner. The resulting sparse set of 3D strokes can be rendered as abstract sketches containing essential 3D characteristic shapes of various objects. We demonstrate that 3Doodle can faithfully express concepts of the original images compared with recent sketch generation approaches. [^1] Code available at https://github.com/changwoonchoi/3Doodle
3D sketch lines, 3D strokes, differentiable rendering
submissionid: 336copyright: rightsretainedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4article: 107publicationmonth: 7doi: 10.1145/3658156ccs: Computing methodologies Parametric curve and surface modelsccs: Computing methodologies Non-photorealistic rendering
@@
Figure 1. We propose 3Doodle, a method to draw 3D strokes from multi-view images. Our approach can express the target object with a highly compact set of geometric primitives, and the recovered 3D strokes are rendered to generate view-coherent sketches.
1. Introduction
@@
Free-hand sketching is an effective tool for visual communication, as only a few sketch lines can summarize perceptually distinctive components. Sketch strokes are also intuitive tools to interact with users in editing an image in many graphics applications (Yu et al., 2019; Zeng et al., 2022). Previous works further explore the connection between 3D structure and conventional sketch lines and show promising results in converting sketches into 3D objects (Li et al., 2020b, 2022; Guillard et al., 2021; Zhang et al., 2021) or extracting sketch lines from detailed 3D geometry (Liu et al., 2020; Bénard et al., 2019; Liu et al., 2021a). Like image editing applications, user sketches provide an interactive tool to edit the 3D models (Mikaeili et al., 2023; Bandyopadhyay et al., 2023).
@@
Sketch primitives with 3D context can be an effective communication tool for both structural and semantic context, widening the scope of potential applications for VR/AR scenarios. The aforementioned works assume full 3D models are available for generating sketches or editing 3D objects with line strokes. However, it is not trivial to formulate the exact mapping from artistic sketch curves on-screen space to the underlying 3D structure of the object. While sketches may contain important spatial arrangements of the object, hand-drawn sketches do not strictly abide by the 3D geometry, either due to a lack of delicate skills or the artist’s creative expression. Therefore, we cannot extract consistent 3D lines from 2D sketch lines extracted individually from different viewpoints. Previous works also propose sophisticated techniques for converting 3D models, such as high-resolution triangular mesh, into sketch images, which result in delicate renderings rather than sparse and abstract strokes (DeCarlo et al., 2003; Judd et al., 2007). The task does not rely on a data-driven method, as it is hard to obtain a dataset of 3D object-sketch pairs in sufficient quantity and quality.
@@
We directly optimize for proposed lightweight 3D primitives given only multi-view images of an object as input. We do not need to individually stylize the input images into sketches, and we also do not require any explicit 3D mesh or NeRF volume (Mildenhall et al., 2021). Most importantly, our rendering pipeline is fully differentiable for a small set of parameters representing the Bézier curves and superquadrics. We can then directly optimize the parameters to minimize the popular perceptual loss functions, such as CLIP (Radford et al., 2021) and LPIPS (Zhang et al., 2018), so that the rendered images are perceived as hand-drawn 2D sketches from various viewpoints. Further, the coherent 3D structure summarizes important geometric features that satisfy multi-view consistency, which can potentially be extended to various 3D manipulation tasks.
@@
We demonstrate that the proposed compact 3D representation can easily mimic an expressive hand drawing of diverse pictured objects, including real-world objects. Our 3Doodle faithfully expresses concepts of the target shape and provides view-consistent sketch images.
2. Related Works
@@
Generating sketches from photos of objects suffers from a large domain gap between pictures and sketches. Images contain dense, detailed evidence of real objects. In contrast, sketches are sparse and abstract, as they are traditionally a quick tool for an artist to summarize the essential components of the given object on an image plane. While extracting edges from image pixels (Canny, 1986; Su et al., 2021) can also create a line drawing of the object, they do not enjoy the compactness and implicit semantic mapping of conventional sketches. Numerous works overcome the domain gap by exploiting datasets composed of pairs of images and sketches (Song et al., 2018; Kampelmuhler and Pinz, 2020). However, obtaining high-quality sketches of given images on a large scale is difficult. Most sketch generation methods only cover specific categories in datasets, such as human portrait (Berger et al., 2013; Yi et al., 2019), or a limited number of object categories  (Yu et al., 2017; Eitz et al., 2012; Ha and Eck, 2017). A few works alleviate the necessity of paired datasets and generate sketches of arbitrary objects with style transfer  (Huang and Belongie, 2017; Liu et al., 2021b) or image translation (Zhu et al., 2017; Isola et al., 2017). Although they can mimic the style of the free-hand sketches to some extent, the generated pixels are still composed of dense information from original images. The images in the converted style do not necessarily maintain compact and sparse representation, which are among the most powerful features of sketches.
3Doodle proposes maintaining the sparse characteristics by directly optimizing a set of strokes, similar to recent approaches (Li et al., 2020a; Vinker et al., 2022). The direct optimization does not require any dataset to train and, therefore, can generate sketches of arbitrary class objects. However, previous works optimize 2D strokes in the input image view without explicit 3D structure. Instead, we optimize 3D strokes with a novel differentiable rendering pipeline. 3Doodle, therefore, generates a compact 3D representation, which can be rendered in various viewpoints to deliver abstract semantic structure.
@@
Non-photorealistic rendering (NPR) renders 3D representations that exhibit various artistic styles. Among possible variations, our sketch generation task particularly relates to line drawing from 3D models. Line drawings reflect how humans understand the world (Bénard et al., 2019). Previous works on line-drawing extraction trace feature lines on surface extrema (Ohtake et al., 2004) in addition to view-dependent geometric boundaries such as occluding contours (Bénard et al., 2019) or suggestive contours (DeCarlo et al., 2003). A more recent method (Liu et al., 2020) combines an entirely geometric approach with image translation and trains a neural network that generates sketches from 3D models. However, all the works rely on having an accurate mesh model, which may not be accessible in real-world scenarios. Further, geometric processing results deviate from typical free-hand drawings because they ignore the effects of texture variations, and lines are densely populated compared to rough sketches.
Recent NPR approaches propose relaxing the mesh requirement and using the 3D volume of neural radiance fields (NeRF) trained from multi-view images instead. They transform pre-trained NeRF into the style of an exemplar image to generate multi-view consistent sketches (Zhang et al., 2022a, 2023). While the pipelines use more casual input than a complete mesh, they need heavy computation first to train a neural volume. Neural Edge Fields (NEF) (Ye et al., 2023) also extracts a compact set of parametric curves as 3Doodle, but it is converted with additional 2D edge input from pre-trained neural fields. Furthermore, the 3D lines from NEF cannot represent view-dependent contours, which are crucial in expressing the overall shape of the object.
3. 3Doodle
Figure 2. Overview of 3Doodle. We generate a compact 3D geometric representation from multi-view images of objects. We separately define view-independent stroke ($\mathcal{S}^{\text{3D}}_{\text{ind}}$) and view-dependent stroke ($\mathcal{S}^{\text{3D}}_{\text{dep}}$). We represent a view-independent stroke as a set of 3D Bézier curves and a view-dependent stroke as a contour of superquadrics. (Sec. 3.1) We also propose a fully differentiable rendering method to render sketches from the 3D strokes. (Sec. 3.2) Finally, our 3D stroke parameters are directly optimized with perceptual losses. (Sec. 3.3)
We first introduce our 3D geometric primitives that generate sketch images in Sec. 3.1. Then our differentiable rendering pipeline in Sec. 3.2 converts the proposed primitives into an abstract sketch image. The compact set of parameters is effectively optimized to minimize the perceptual loss function as described in Sec. 3.3.
3.1. Sketch Representation
@@
(1) $\mathcal{S}^{\text{3D}}=\mathcal{S}_{\text{ind}}^{\text{3D}}\cup\mathcal{S}_{\text{dep}}^{\text{3D}}.$
@@
(2) $\mathcal{S}_{\text{ind}}^{3\text{D}}=\left\{B^{\text{3D}}(p_{i})\right\}_{i=1}^{N_{\text{ind}}},\quad p_{i}=(p_{i}^{0},p_{i}^{1},p_{i}^{2},p_{i}^{3})$
@@
(3) $B^{\text{3D}}(t;p_{i})=\sum_{j=0}^{3}b_{j}(t)p_{i}^{j},\quad b_{j}(t)={\binom{3}{j}}t^{j}(1-t)^{3-j},$
@@
Additionally, view-dependent components $\mathcal{S}_{\text{dep}}^{3\text{D}}$ encapsulate the 3D volume of the given object. Inspired by the previous works (Pentland, 1986; Alaniz et al., 2023; Paschalidou et al., 2019, 2020), we use the composition of superquadradics (Barr, 1981) as a compact parametric representation that can express various 3D shapes. More concretely,
(4) $\mathcal{S}_{\text{dep}}^{3\text{D}}(\mathbf{d})=\sigma_{\text{contour}}\left(\bigcup_{i=1}^{N_{\text{dep}}}S(\theta_{i}),\mathbf{d}\right),$
where $N_{\text{dep}}$ is the number of superquadrics, and $\sigma_{\text{contour}}(G,\mathbf{d})$ is the view-dependent volume density function that generates the contour of any geometric volume $G$ at the given viewing direction $\mathbf{d}$. We will discuss contour extraction in Sec. 3.2.
@@
(5) $f(\mathbf{x};\alpha_{i},\epsilon_{i})=\left(\left(\frac{x}{\alpha_{i,1}}\right)^{\frac{2}{\epsilon_{i,2}}}+\left(\frac{y}{\alpha_{i,2}}\right)^{\frac{2}{\epsilon_{i,2}}}\right)^{\frac{\epsilon_{i,2}}{\epsilon_{i,1}}}+\left(\frac{z}{\alpha_{i,3}}\right)^{\frac{2}{\epsilon_{i,1}}},$
@@
(6) $S(\mathbf{x};\theta_{i})=f\left(\mathbf{R}_{i}^{-1}(\mathbf{x}-\mathbf{t}_{i});\alpha_{i},\epsilon_{i}\right).$
@@
(7) $S(\mathbf{x};\theta)=\bigcup_{i=1}^{N_{\text{dep}}}S(\mathbf{x};\theta_{i})=\min\limits_{i}S(\mathbf{x};\theta_{i}).$
3.2. Differentiable Rendering of Sketch Lines
Given the 3D geometric primitives, we can draw black sketch lines on a white canvas $\mathcal{R}(\mathcal{S}^{\text{3D}})$. A complete sketch is the union of view-independent and view-dependent components. Similar to differentiable rendering of parametric objects (Worchel and Alexa, 2023), we propose a fully differentiable rendering pipeline to directly optimize the 3D parameters to produce abstract sketch images of the target objects.
@@ 3.2. Differentiable Rendering of Sketch Lines
We provide the proof of Theorem 1 in Appendix A. Following Theorem 1, we can render the 3D Bézier curve $B^{\text{3D}}$ at a given viewpoint by 1) perspective-projecting four control points and then 2) rendering a 2D Bézier curve with existing differentiable rasterizer (Li et al., 2020a). The rendering pipeline is fully differentiable since the projection operation and rasterizer are both differentiable.
Figure 3. (a) We visualize the volume density $\sigma_{\text{vol}}$ and surface volume density $\sigma_{\text{surf}}$. (b) We display the volume-rendered results of each volume density component. One can obtain the contour of geometric volume by volume rendering our proposed view-dependent contour volume density $\sigma_{\text{contour}}$.
@@
We provide the formulation to generate contour sketches that are differentiable to the superquadric parameters. Volume rendering exploits smooth transition in volume density and provides a powerful tool to propagate image-space gradients to 3D geometry, as demonstrated by neural radiance fields (Mildenhall et al., 2021). Similarly, we propose a differentiable pipeline by defining smooth density fields representing sketch lines. To allow continuous variation, ISCO (Alaniz et al., 2023) proposes volume density of superquadric
(8) $\sigma_{\text{vol}}(\mathbf{x})=\text{sigmoid}(\gamma(1-S(\mathbf{x})),$
where $S$ is an implicit surface function defined in Eqs. 5, 6 and 7. $\gamma$ controls the slope of the shape boundary. For sketch generation, we revise the volume density function to be concentrated near the surface $S=1$:
(9) $\sigma_{\text{surf}}(\mathbf{x})=a\cdot\text{sigmoid}\left(\frac{1}{\gamma(1-S(\mathbf{x}))^{2}+\epsilon}-\gamma(1-S(\mathbf{x}))^{2}-b\right),$
where $a,b\in\mathbb{R}^{+}$ are hyperparameters determining the intensity and thickness of the surface volume. We illustrate the $\sigma_{\text{vol}}$ and $\sigma_{\text{surf}}$ in one-dimensional example in Fig. 3 (a).
@@
(10) $\sigma_{\text{contour}}(\mathbf{x},\mathbf{d})=(1-(\mathbf{n}(\mathbf{x})\cdot\mathbf{d})^{\beta})\sigma_{\text{surf}},$
where $\mathbf{d}$ is a viewing direction and $\mathbf{n}$ is a normal direction at the querying point $\mathbf{x}$. Note that the normal direction $\mathbf{n}(\mathbf{x})=\left[\frac{\partial S}{\partial x},\frac{\partial S}{\partial y},\frac{\partial S}{\partial z}\right]$ can be calculated in a closed form with analytic derivative. Then we render the view-dependent sketch $\mathcal{R}(\mathcal{S}_{\text{dep}}^{\text{{3D}}})$ using volume rendering technique and numerical quadrature approximation (Max, 1995) with our contour volume density $\sigma_{\text{contour}}$.
3.3. Optimization
Given the differentiable pipeline, we optimize the parameters of our sparse 3D primitives to generate sketch images. While there exists a significant domain gap between photo-realistic input images and the abstract sketch (Vinker et al., 2022), we design a loss function to encourage a balance between structural components and semantic perception compared to input multi-view images. Specifically, we use LPIPS loss (Zhang et al., 2018) to capture rough geometric layout and CLIP score (Radford et al., 2021) to maintain the high-level semantic meaning:
(11) $\begin{split}\mathcal{L}=\sum_{I\in\mathcal{I}}\lambda~{}\rho\left(\text{LPIPS}(I,\mathcal{R}(\mathcal{S}^{\text{3D}})),\alpha,c\right)\\ +\,\text{dist}(\text{CLIP}(I),\text{CLIP}(\mathcal{R}(\mathcal{S}^{\text{3D}}))),\end{split}$
where $I$ is the image sampled from the input multi-view images $\mathcal{I}$, and $\text{dist}(a,b)=1-\frac{a\cdot b}{\|a\|\cdot\|b\|}$ is a cosine distance. $\rho(x,\alpha,c)$ is a robust loss function (Barron, 2019), which stabilize the optimization despite outliers. Outliers are inevitable as it is not trivial to correctly account for view-dependent occlusion only with line primitives. We use $\alpha=1,c=0.1$ for all experiments. More studies regarding the effects of loss terms can be found in the Sec. 4. Our sketch parameters are directly optimized with Adam optimizer (Kingma and Ba, 2014) by minimizing the loss term in Eq. 11.
Due to the highly non-convex nature of our minimizing objective, appropriate initialization leads to fast and robust optimization. We leverage the SfM (structure-from-motion) point clouds obtained while estimating the camera poses from input multi-view images. We sample SfM points using farthest point sampling and initialize them to be locations of 3D Bézier curve control points and the center of superquadrics. We further discuss the effects of the initialization method in Sec. 4.
4. Experiments